OpenStudy (anonymous):

@terenzreignz Could you help me with this one too? =) A meteorologist reported an 82% chance the weather would be sunny and a 74% chance it would be sunny and that the temperature would get over 85 °F. If it is a sunny day, what is the probability the temperature will also be more than 85 °F? a. 31.7% b. 74.0% c. 90.2% d. 18.0%

terenzreignz (terenzreignz):

This is a little tricky. We need what's called 'conditional' probabilities... \[\Large Pr(A|B) = \frac{Pr(A \ and \ B)}{Pr(B)}\]

terenzreignz (terenzreignz):

To translate... Probability of event A happening given that event B happens is equal to the probability of both A and B happening divided by the probability of B happening.

terenzreignz (terenzreignz):

So here, our event A is that the temperature would go over 85 degrees. Our event B is that it would be sunny. So, (A|B) is the probability of the temperature going over 85 degrees given that it is sunny. Okay?

OpenStudy (anonymous):

Got it!

terenzreignz (terenzreignz):

So, what's the probability that both A and B occur? (IE, that it is both sunny and the temperature goes over 85?)

OpenStudy (nincompoop):

she said she got it yo

terenzreignz (terenzreignz):

I suppose.

OpenStudy (anonymous):

I think the answer is C. =)

terenzreignz (terenzreignz):

Well, who said it wasn't? :P \[\Large Pr(A \ and \ B) = 0.74\]\[\Large Pr(B) = 0.82\] \[\Large Pr(A|B) = \frac{Pr(A \ and \ B)}{Pr(B)}=\frac{0.74}{0.82}\approx0.9024 \approx90.2\%\]

OpenStudy (anonymous):

Thank you very much for your help!

OpenStudy (nincompoop):

thank you for answering her problem, @terenzreignz